It is estimated that t years from now the value V (in dollars) of an acre of land near the ghost town of Cherokee, California, will be increasing at the rate of V(_^') (t) =〖8t〗^3/√(〖0.2t〗^4+8000) dollars per year. If the land is currently worth $500 per acre, how much will it be worth in 10 years?
Answer #1

we are given

V'(t)=\frac{8t^3}{\sqrt{(0.2t)^4+8000}}

we can integrate it

V(t)=\int \frac{8t^3}{\sqrt{(0.2t)^4+8000}}dt

we can use u-subs

u=(0.2t)^4+8000

u=0.0016\times 4t^3dt

u=0.0064ť dt

V(t)=\int \frac{8t^3}{\sqrt{u}}\times \frac{du}{0.0064t^3}

V(t)=\int \frac{8}{\sqrt{u}}\times \frac{du}{0.0064}

=1250\times \int \frac{1}{\sqrt{u}}du

=1250\times \int \:u^{-\frac{1}{2}}du

=2500\sqrt{u}+C

now, we can plug back u

and we get

V(t)=2500\sqrt{(0.2t)^4+8000}+C

we are given

at t=0, V=500

and then we can solve for C

500=2500\sqrt{(0.2(0))^4+8000}+C

C=-223106.80

now, we can plug back C

V(t)=2500\sqrt{(0.2t)^4+8000}-223106.80

now, we can plug t=10

V(10)=2500\sqrt{(0.2(10))^4+8000}-223106.80

V(10)=723.492.........Answer